### v1

# How Linguistics Learned to Stop Worrying and Love the Language Models

Abstract
Language models can produce fluent, grammatical text. Nonetheless, some maintain that language models don’t really learn language and also that, even if they did, that would not be informative for the study of human learning and processing. On the other side, there have been claims that the success of LMs obviates the need for studying linguistic theory and structure. We argue that both extremes are wrong. LMs can contribute to fundamental questions about linguistic structure, language processing, and learning. They force us to rethink arguments about learning and are informative for major questions in linguistic theory. But they do not replace linguistic structure and theory. We offer an optimistic take on the relationship between language models and linguistics.
Richard Futrell
University of California Irvine, USA
rfutrell@uci.edu
Kyle Mahowald
The University of Texas at Austin, USA
kyle@utexas.edu
Keywords: language models, linguistic theory, language learning, information theory, functional linguistics, statistical learning, neural networks
1 Introduction
It’s 1968, and Norm and Claudette are having lunch. Norm is explaining his position that all human languages share deep underlying structure and has worked out careful theories showing how the surface forms of language can be derived from these underlying principles. Claudette, whose favorite movie is the recently released 2001: A Space Odyssey and who particularly loves the HAL character, wants to make machines that could talk with us in any human language. Claudette asks Norm whether Norm thinks his theories could be useful for building such a system. Norm says he is interested in human language and the human mind, found HAL creepy, and isn’t sure why Claudette is so interested in building chatbots or what good would come of that. Nonetheless, they both agree that it seems likely that, if Norm’s theories are right (and he sure thinks they are!), they could be used to work out the fundamental rules and operations underlying human language in general—and that should, in principle, prove useful for building Claudette’s linguistic machines. Claudette is very open to this possibility: all she wants is a machine that talks and understands. She doesn’t really care how it happens. Norm and Claudette have very different goals, but they enjoy their conversations and are optimistic that they can both help each other.
Fast forward to 2025. Norm has worked for decades on a variety of diverse languages, developing explanatory theories of linguistic structure. Claudette got more and more interested in engineering, amassing huge amounts of data, and training statistical models. Norm and Claudette slowly fell out of touch over the years and haven’t seen each other for a while. They are planning another lunch: what should they talk about?
Just like Claudette always dreamed of, she now has machines that appear to be able to talk with her in English, producing grammatical and seemingly sensible utterances. How relevant is it that the architecture of Claudette’s machines seems to have nothing to do with the structure of language as identified by Norm and other linguists and cognitive scientists? Claudette might not care: she just wanted a system that worked and this stuff works. On the other hand, what if Norm was right about the nature of language—does that mean the machines aren’t actually as impressive as Claudette thinks, because they are relying on shallow pattern matching? Or are Claudette’s machines evidence that Norm’s theories were wrong? More generally, what insights do these machines have for the scientific study of language?
One approach Norm could take would be to dismiss Claudette’s machines out of hand as irrelevant to human language, as they are not realistic models of the human brain, nor are they intended to be. Under this view, neural networks are no more relevant to linguistics than submarine engineering is to an ichthyologist—just because both submarines and fish can move underwater does not mean that you can learn much about one from studying the other.
This negative view takes several forms. One argument is that Language Models (yes, Claudette’s machines are Language Models; henceforth LMs) are so fundamentally different in their architecture from people, and have access to so much more data, that whatever they are doing is so different as to be totally irrelevant for humans (Bolhuis et al.,, 2024; Chomsky et al.,, 2023). Another argument is that neural sequence architectures could learn anything, so the fact that they learn language (to the extent that they do) is uninformative (Rawski and Baumont,, 2023; Moro et al.,, 2023; Chomsky,, 2023; Chomsky et al.,, 2023; Bolhuis et al.,, 2024). Under this view, neural networks are like mathematical clay that can be formed into any shape (Robinson,, 2023): the fact that they can be formed into something like the shape of language doesn’t tell us anything about language, any more than the fact that we can form clay into the shape of a human body tells us anything about anatomy. A distinct but related negative approach has been to deny that language models have learned, or ever could learn, the putatively key properties of human language, and thus are either irrelevant to language science or do not challenge any existing theories (Lan et al.,, 2024; Fox and Katzir,, 2024). We believe these negative views are short-sighted: language models do learn nontrivial aspects of linguistic structure, and they do give important insights that change how we should think about language. As language scientists, we ignore them at our peril.
An opposite approach is to dismiss traditional linguistic ideas of structure, on the grounds that they turned out to be either useless or of negative value in engineered language technologies (Piantadosi,, 2023). Under this view, decades of work on the structure of language turned out to be barking up the wrong tree, and the theory of language should be rebuilt on top of the only computational foundation that has produced a machine that can actually use language: large neural networks. We believe this view is also short-sighted. It throws out hard-won analytical discoveries about the structure of language and leaves us adrift in terms of a scientific theory of language, without a clear way to approach the question of why human language is the way it is, or even a sense of what the interesting questions are. Moreover, language models are currently most successful in languages with internet-scale data and are biased towards standardized prestige dialects, neglecting under-resourced languages and dialects (Blasi et al.,, 2022). Work using language models will benefit from linguistic work focused on communities outside this current narrow band—which requires the expertise of documentary linguists, sociolinguists, anthropologists, and community stakeholders.
Thus, we advocate a third view (joining others in cognitive science and philosophy who have advocated for this middle ground, for example, Smolensky,, 1988; Millière,, 2024; Portelance and Jasbi,, 2023; Pater,, 2019; Potts,, 2025): large language models are not a complete theory of language—in fact, no one has such a theory—but they are hugely informative about language and its structure, learning, processing, and relationship with the larger structure of the mind. Language models have set off an intellectual explosion in cognitive science, machine learning, philosophy of mind, and other fields, in which longstanding ideas have been overturned, truly novel ideas are emerging, and disciplinary boundaries are dissolving. Linguistics has a chance to stand at the center of this huge intellectual foment, and would be remiss to isolate itself intellectually on the basis that language models don’t look like any particular existing theory. We believe that language science can contribute (and in fact already has contributed) to the development of language models, and language models can contribute (and in fact already have contributed) insights about language. Norm and Claudette have a lot to talk about.
2 Statistical models of language have outperformed expectations
2.1 A brief history of statistical language learning
The generative school of modern linguistics arose from an argument against the idea that the structure of language could be learned from the statistics of language data. Chomsky, perhaps most famously illustrated this by contrasting “(1) Colorless green ideas sleep furiously” with “(2) Furiously sleep ideas green colorless.” Even though neither sentence makes semantic sense, (1) is clearly grammatical in a way that (2) isn’t. As Chomsky, (1957) put it: “The notion ‘grammatical in English’ cannot be identified in any way with the notion ‘high order of statistical approximation to English.’ It is fair to assume that neither sentence (1) nor (2) (nor indeed any part of these sentences) has ever occurred in an English discourse. Hence, in any statistical model for grammaticalness, these sentences will be ruled out on identical grounds as equally ‘remote’ from English.”
The conclusion from these arguments was that linguistic structure could only be characterized in terms of formal systems, based on rules or constraints and operating over structured arrays of symbols (Chomsky,, 1965). Early on, the expectation was that such systems would form the basis for language technologies such as machine translation (Hays,, 1960; Hutchins,, 1979). It was believed that these formal systems should be constructed by linguists since the task of learning them from data was hard or impossible. Yet despite myriad efforts to build machine translation systems, grammatical parsers, and other tools, linguistic competence remained elusive for machines. Symbolic approaches that sought to elucidate rules and structures often proved unable to capture all the exceptions and complexity that characterized natural language as it is used.
By the late 1980s and 1990s, statistical learning had a major renaissance in natural language processing (Brown et al.,, 1990; Manning and Schütze,, 1999; Pereira,, 2000), and started to reappear in the human language learning literature as well (Saffran et al.,, 1996). Yet despite success at various NLP tasks, these techniques seemed frustratingly unable to get past approaches that simply counted up words and phrases (Wang and Manning,, 2012; Arora et al.,, 2017). The connectionist movement in the 1980s and 90s seemed promising (Rumelhart and McClelland, 1986b, ; Smolensky,, 1988; Elman, 1990b, ) (and turned out to be the most important forerunner of modern language models), but there were fundamental concerns about the ability of these approaches to scale up and to represent the rules and structures that seem to characterize language (Pinker and Prince,, 1988), despite accounts of how connectionist models could in principle implement rule-like symbolic behavior (Smolensky,, 1990).
Even researchers optimistic about the role of statistical learning in language acquisition were, in the recent past, deeply skeptical that end-to-end neural approaches would succeed on interesting linguistic tasks. Representatively, Chater et al., (2006) wrote, “connectionism is no panacea here; indeed, connectionist simulations of language learning typically use small artificial languages, and, despite having considerable psychological interest, they often scale poorly.” Tenenbaum et al., (2011) wrote, “connectionist models sidestep these challenges by denying that brains actually encode such rich knowledge, but this runs counter to the strong consensus in cognitive science and artificial intelligence that symbols and structures are essential for thought.” Nevertheless, given the way that linguistics and machine learning developed and evolved, it now sometimes seems to be taken for granted in practice that ideas from linguistic theory will not form the basis of proficient language processing systems.
Throughout the statistical renaissance, generative linguists continued to claim that statistical methods would never solve interesting problems related to learning linguistic structure. In a review article, Everaert et al., (2015) reiterated the claim that statistical approaches would be fundamentally unable to distinguish sentences like “How many cars did they ask if the mechanics fixed?” from ungrammatical ones like “*How many mechanics did they ask if fixed the cars?”, or that they would only be able to do so if provided with certain built-in formal structures. Berwick et al., (2011) were skeptical that recurrent neural networks could ever be much more powerful than bigram models.
Against this backdrop, the fortunes of connectionist models started changing in the 2010s—as a result of new techniques and, most importantly, due to increased computational power that enabled training neural models much more efficiently (Hinton et al.,, 2006). By 2016, neural models showed rudiments of structure-dependent generalizations like subject–verb agreement (Linzen et al.,, 2016). Over the coming years, the success of models at acquiring structure-dependent linguistic abilities continued to grow (Futrell et al., 2019a, ; Wilcox et al.,, 2018; Manning et al.,, 2020; Hu et al.,, 2020; Warstadt and Bowman,, 2022; Mahowald et al.,, 2024).
The growth from early neural models in, e.g., 2011 to now is, from a historical perspective, remarkable and surprising to virtually everyone in the field. Sutskever et al., (2011) introduced an at-the-time state-of-the-art recurrent neural network that produced output like “In the show’s agreement unanimously resurfaced. The wild pasteured with consistent street forests were incorporated by the 15th century BE.” and praised it as a “surprisingly good language model,” noting the “richness of their vocabularies,” that the “text is mostly grammatical,” and “parentheses are usually balanced over many characters.” Compared to today’s language models, the consistent street forests seem a long way away: 2011 might as well have been the 15th century BE. Models like ChatGPT or Claude, based on large neural networks, produce seemingly endless streams of highly fluent and entirely grammatical (if not always perfectly humanlike) text, and form the basis of the most successful NLP approaches.
2.2 Neural LMs learn nontrivial linguistic structure
The rapid development of these capacities has been treated with healthy skepticism. The field of linguistics has provided a deep body of work exploring the subtleties and complexities of even seemingly simple linguistic utterances, and LMs have access to vast amounts of data. There is good reason to be skeptical and wonder whether the models are simply memorizing, or giving the appearance of linguistic competence without really having it.
The most naïve way that one might think to test for grammatical competence in language models would be to see whether they assign higher probability to grammatical strings than ungrammatical ones. However, this would not work, even for an ideal statistical model, for reasons related to the point that Chomsky, (1957) was making. For example, the ungrammatical string ‘Snails died the old’ has a probability of $\sim 2^{-49}$ under the GPT-2 language model, whereas the grammatical string ‘The ancient crustaceans expired’ has a lower probability of $\sim 2^{-55}$, simply because it uses lower-frequency words (Wilcox et al., 2023a, ).
But the failure of this naïve approach does not mean that langauge models lack grammatical knowledge. In general—for language models and in real language use—there are many factors that influence the probability of a string, of which the grammar of the language (as represented in the language model or in the mind of a human speaker) is only one: word frequency, utterance length, online processing constraints such as memory limitations, and plausibility given world knowledge all feed into the probability of an utterance. Indeed, not only utterance probabilities, but also comprehension accuracy, reaction time, and indeed any psychometric dependent variable are affected by all of these factors jointly—including the subjective grammaticality judgments that form the basis of much of the formal syntax literature (Kluender and Kutas,, 1993; Hofmeister et al.,, 2013; Mahowald et al.,, 2016; Lau et al.,, 2017). Therefore, if we want to look for evidence that language models represent linguistic structure, we need to search more carefully. [^2] Nor is it sensible to evaluate models based on simply prompting them and asking whether sentences are grammatical or not (Hu and Levy,, 2023) (as done by Dentella et al.,, 2023). To see this, consider that many linguists have held that grammatical knowledge can be represented by a formal system like a context-free grammar; even if a context-free grammar is a good model of linguistic knowledge, it could not be prompted in natural language to respond about whether a sentence is grammatical or not. The ability to do explicit grammaticality judgments when prompted in natural language is not a necessary consequence of grammatical knowledge.
The key is to isolate linguistic structure from these other factors through controlled experimental studies and through probing and manipulating LMs’ internals. Experimentally, from sufficient performance data, one may infer an underlying formal cognitive structure, no matter whether the implementation substrate is a brain or a neural network (Piantadosi and Gallistel,, 2024). This is the standard procedure in linguistics, where data consisting primarily of acceptability judgments is used to postulate underlying linguistic competence. This approach can be applied just as well to LMs.
One form of such a study is to conduct behavioral comparisons of minimally different sentences. For example, LMs may be fed sentences such as The keys to the cabinet are on the table and the ungrammatical The keys to the cabinet is on the table, and the conditional probabilities assigned to the verb form are versus is are compared: if the model understands how agreement works in English, then the probability for are should be higher than is. This comparison controls for plausibility, since the context is the same for the two alternatives. The task could not be solved by any $n$-gram model with $n<5$ (and higher-order $n$-gram models can be ruled out by adding more words between keys and is/are in the stimuli). The lexical frequency of is versus are can be controlled through a more elaborate experimental design, with four conditions in a $2\times 2$ design crossing the grammatical number of the subject with the grammatical form of the verb (as done by Marvin and Linzen,, 2018). These are usual procedures in psycholinguistics, where experimenters develop controlled sets of sentences and then measure dependent variables like reading times. The same methodology can be applied to language models with probability as the dependent variable, aiming to ascertain whether models are sensitive to linguistic structure (Linzen et al.,, 2016; Futrell et al., 2019b, ).
Such studies have revealed behavioral patterns consistent with neural networks representing formal linguistic structures, in cases such as subject–verb agreement (Linzen et al.,, 2016; Bernardy and Lappin,, 2017; Gulordava et al.,, 2018), filler–gap dependencies (Wilcox et al.,, 2018; Wilcox et al., 2023a, ; Kobzeva et al.,, 2023; Suijkerbuijk et al.,, 2023), and recursive embedding of clauses (Futrell et al., 2019b, ; Wilcox et al., 2019a, ; Hu et al.,, 2020), all of which involve highly nontrivial formal structures which statistical models failed to capture in previous work. Example results for subject–verb agreement from GPT-2 are shown in Figure 1: we see that grammatical verb forms are relatively more probable than matched ungrammatical verb forms in all but a few cases (human accuracy in producing the right verb forms in such sentences is 85%; Marvin and Linzen,, 2018). [^3] It has been objected that, while humans make errors in verb agreement and other syntactic constructions, they have an underlying error-free competence which may be distinguished from their errorful performance (Chomsky,, 1965, Ch. 1), whereas the neural networks only have errorful performance with no underlying competence (Fox and Katzir,, 2024). We believe this argument is mistaken for two reasons. First, it is indeed possible to define (multiple) competence–performance distinction(s) in neural networks, for example by separating the knowledge contained in embeddings versus the performance of a linear decoder in extracting that knowledge (Pimentel et al.,, 2020; White et al.,, 2021; Csordás et al.,, 2024). Second, even for humans, we only ever have access to performance data, and we infer mental constructs such as competence from it in order to explain that data (Piantadosi and Gallistel,, 2024). We may make the same inferences based on neural network performance data. Competence is not a datum to be explained, but rather a theoretical construct that is useful for explaining performance data—it is a real pattern (Dennett,, 1989; Nefdt,, 2023) in the sense we will discuss in Section 4.1. The results indicate that the model can represent the non-local structural dependency between the subject of the sentence and the matrix verb.
[…]
A. Matching context, varying verb
B. Matching verb, varying context
Figure 1: Language models show sensitivity to linguistic structure. Figure shows data from targeted evaluation of subject–verb agreement in GPT-2 from Gauthier et al., (2020), based on design and materials from Marvin and Linzen, (2018). Each point represents the log conditional probability for a verb (e.g., singular is or plural are) in a particular sentence. A. Blue points are for grammatical verbs, and red for ungrammatical, in matching contexts. The difference in log probability between grammatical and ungrammatical verbs in identical contexts is indicated by a line. For plural subjects, the grammatical verb is always higher probability. For singular subjects, the grammatical verb is higher probability in 79% of cases. Human accuracy on this task is 85% (Marvin and Linzen,, 2018). B. Blue points for grammatical verbs, and red points for the same verbs in a context that makes them ungrammatical. The correct context makes the grammatical verb higher probability in all cases for the singular verb is, and in 95% of cases for the plural verb are.
Of course, this work has not all been positive in its results (for example, Weissweiler et al., 2023b, ), and it is possible for models to be “right for the wrong reasons” (McCoy et al.,, 2019), adopting shallow heuristics which make correct predictions on certain test sets but do not generalize properly, or models may show mastery of linguistic form without a concomitant ability to understand the implications of utterances (Weissweiler et al.,, 2022)—in effect mastering linguistic form without mastering linguistic function (Mahowald et al.,, 2024). As an example of how shallow heuristics may lead to the illusion of good performance, Vázquez Martínez et al., (2023) find that a simple $n$-gram model performs surprisingly well on the paired grammatical and ungrammatical sentences of the BLiMP dataset (Warstadt et al.,, 2020), which turn out to be poorly controlled for $n$-gram frequency. [^4] We note that this criticism does not apply to studies based on more carefully selected sets of sentences (for example, Marvin and Linzen,, 2018; Futrell et al., 2019b, ; Hu et al.,, 2020; Wilcox et al., 2023a, ).
The possibility that models generalize on the basis of shallow heuristics motivates deeper investigation. We mention two approaches here. The first involves controlling the training data for language models and then observing their generalizations on evaluation data that is unlike anything in their training data. This approach allows one to study true generalization in language models, rather than shallow memorized patterns. For instance, Misra and Mahowald, (2024) found that the relatively rare A+Adjective+Numeral+Noun construction “a beautiful five days in Texas” is learned to some extent by models in which instances of the construction are systematically withheld during training—suggesting generalization. Ahuja et al., (2024), building on McCoy et al., (2020), train Transformer language models on a corpus of English-like text that has been constrained so that subjects and verbs are always adjacent. That is, the corpus contains sentences like I saw the key to the cabinets, but never something like The key to the cabinets is on the table. The question is, then, if the model trained in this way will generalize in the way that humans do, preferring The key to the cabinets is on the table (where the form of the verb depends on the subject of the sentence) over something like The key to the cabinets are on the table (where the form of the verb depends on the linearly previous verb, which is perfectly consistent with the training data). Ahuja et al., (2024) find that Transformer language models do make the human-like generalization, showing that they can learn the formal structure underlying nonlocal subject–verb agreement without ever observing nonlocal subject–verb agreement (see also Jumelet and Zuidema,, 2023; Patil et al.,, 2024).
Another approach is to dig into the language models’ internals, using approaches such as probing, where one attempts to decode linguistic representations from the internal representations of neural networks, and causal interventions, where model internals are changed and the resulting output changes are observed. Such studies have increasingly revealed genuine representations of grammatical structure once neural networks’ internal representations are decoded properly (Lakretz et al.,, 2019; Voita and Titov,, 2020; Manning et al.,, 2020; Papadimitriou et al.,, 2021; Ravfogel et al.,, 2021; Lampinen,, 2024; Diego-Simón et al.,, 2024, among others). For example, Hewitt and Manning, (2019) and Chi et al., (2020) find that syntactic relationships among words in sentences are encoded geometrically in a linear subspace of the high-dimensional contextual embedding space for words. Ravfogel et al., (2021) showed that these syntactic representations are causally active: systematically altering them changes model behavior in the ways we would expect.
An important caveat to these findings is that the various evaluations of structural representation in language models have mostly used English as the target language (Blasi et al.,, 2022). As such, current architectures and paradigms are likely biased towards English and other widely spoken languages. An important area for future work is to extend the empirical scope of these studies to other languages whose formal patterns may differ from English (Bender,, 2022).
There are ongoing debates about the extent to which language models capture more complex formal patterns within languages (Vázquez Martínez et al.,, 2023; Lan et al.,, 2024) or to what extent they can be said to “understand” (Bender and Koller,, 2020) or refer to things in the world (Mandelkern and Linzen,, 2023; Lederman and Mahowald,, 2024). This is not the right place to adjudicate these debates. What we think is uncontroversial, based on the work summarized above, is that language models have learned nontrivial formal patterns corresponding to linguistic structure, to a greater extent than many expected was possible. Which is to say, we believe that even if language models are not perfect at grammar, they nevertheless really have learned something about linguistic structure, in a way that problematizes earlier claims about the role of statistics in language learning. They are not purely relying on heuristics or memorization or cheap tricks. They have, to a meaningful extent, learned “the real thing”—that is, the thing that we care about, as language scientists interested in questions like how languages are learned, how they are processed, how and why they vary, and where they come from.
We now proceed to ask why this is interesting for the science of language, and how the science of language productively integrate these insights.
3 The success of LMs is interesting for the science of language
“…nor did Pnin, as a teacher, ever presume to approach the loft halls of modern scientific linguistics, that ascetic fraternity of phonemes, that temple wherein young people are taught not the language itself, but the method of teaching others to teach that method; which method, like a waterfall splashing from rock to rock, ceases to be a medium of rational navigation but perhaps in some fabulous future may become instrumental in evolving esoteric dialects—Basic Basque and so forth—spoken only by certain elaborate machines.”
Vladimir Nabokov
Pnin
LMs learn nontrivial aspects of language. But one could still object that this result is not relevant to language science because human language is the product of the human brain (and/or vocal tract, body, social environment, etc.), whereas LMs have a totally different architecture and different objective. Perhaps these differences are so great that, just as there is limited value to the ichthyologist studying a submarine, a language scientist has nothing to gain from studying language models.
Here we argue that the success of language models is relevant, indeed important, to the science of language, because (1) there are good reasons to believe that there should be parallels between engineering artifacts like language models and human language, and (2) the success of language models in learning from text upends ways of thinking that are deeply ingrained in generative linguistics and parts of cognitive science, and (3) language models are already connected to statistical and probabilistic traditions in linguistics, and in fact arose out of them to a large extent. While some ascribe the success of LMs to engineering alone, LMs’ success is also a success of an intellectual tradition within linguistics. As such, LMs aren’t alien invaders into linguistics from engineering. Rather, they are tools of the kind that have long been used to answer fundamentally linguistic questions.
3.1 Parallels between engineering models and cognition
From the study of vision, there is strong precedent for the idea that neural networks developed purely for practical applications can tell us a great deal about cognition as it is implemented in the brain. Similarly, we think there is good reason to think that LMs will be informative for the study of human language.
Hubel and Wiesel, (1959) discovered that early processing of visual information is performed by neurons that are selectively responsive to edges. After this discovery, the question became why early visual processing works this way. Olshausen and Field, (1996) showed that edge filters resembling those found in the visual cortex were the generic solution to the problem of representing visual information accurately in a neural network under a constraint that only a small number of units should be active (that is, a sparsity constraint, arising ultimately from a power constraint on neural firing). Even better explanation came from the engineering of artificial visual systems, in particular the development of AlexNet, a large (for the time) hierarchical convolutional neural network for image classification (Krizhevsky et al.,, 2012). Yamins et al., (2014) showed that this architecture, when trained to do object recognition using the same engineering methods and techniques as in the machine learning model, not only developed edge filters in its early layers, but also receptive fields in later layers corresponding to later layers of visual processing in the primate brain. The overall picture that emerges from this line of work is: the neuronal organization of visual processing is determined by the function of finding a sparse code to identify and manipulate objects in the environment, a function largely shared between the biological and artificial systems.
Here, the artificial system provided guiding insights for understanding the natural system. Something similar is happening in the science of language processing, where the internal representations developed by language models are predictive of activation patterns in the human language network (Goldstein et al.,, 2022; Rathi et al.,, 2024). The picture is not yet as clear as it is in vision—this is an area of active research, and there are no animal models we can use to understand language in the brain—but there is precedent to think that the artificial system will be informative about the natural system here too.
Why did this work in vision? To explain the success of the approach, Cao and Yamins, (2021) introduce the Contravariance Principle: in short, if a computational problem is hard, then there are only a few ways to solve it, so we expect different systems (for example, neural networks and the brain) to converge to the same solution (see also Huh et al.,, 2024). Vision is hard, therefore we can expect that any system that solves the problem will share core features with any other. We see this not only in the comparison of humans vs. neural networks, but also the comparison of humans vs. monkeys (Rajalingham et al.,, 2018), and primates vs. animals whose visual cortices arise from totally different evolutionary phylogenies, such as cephalopods (Pungor et al.,, 2023).
It seems undeniable that learning and using language is hard, and as we have discussed, large Transformer-based models do learn and deploy nontrivial aspects of language. If the Contravariance Principle is right, then it is reasonable to expect some degree of commonality in the solutions that exist in the brain and in Transformers. Indeed, the massive literature on linguistic interpretability in Transformers, which we partially reviewed in Section 2.2, has revealed representational strategies in neural networks that lead to correct novel predictions about human performance (Lakretz et al.,, 2021), and the way that Transformers process syntactic features like agreement cues has intriguing parallels with independently-proposed cognitive frameworks for modeling human language processing (Lewis and Vasishth,, 2005; Ryu and Lewis,, 2021; Timkey and Linzen,, 2023).
It is just not the case that these models can be dismissed out of hand because they are different from the brain. To do so would be to miss out on a source of in silico insight and inspiration which has been hugely useful in other areas of cognition.
3.2 Understanding LM success requires rethinking language learning
3.2.1 The significance of the learning problem in linguistics and cognitive science
For a century, cognitive scientists, linguists, computer scientists, and philosophers developed formal theories of how language learning—or any learning—must work. The key concept, known under many names, has been inductive bias (Goyal and Bengio,, 2022), as illustrated in a few ways in Figure 3.2.1. We can think of a learner as a device that takes in some data and outputs a hypothesis, or a set of hypotheses, or a probability distribution over hypotheses, for the underlying process generating the data. For example, given a bunch of data points in a 2D plane, one hypothesis would be that the points are generated along a line in the plane and then perturbed by noise; this is the assumption underlying linear regression. Given linguistic utterances as data, one might hypothesize a certain grammar underlying the data.
The key issue is that, in a very general sense, the data fundamentally underdetermine the hypothesis that the learner arrives at, because any data is compatible with many, many hypotheses. In Figure 3.2.1A, where the learning data is points and the hypotheses are lines, both the straight and the curvy lines fit the learning data equally well; the fact that the straight line seems like the right hypothesis—and is in fact the one that humans would seize on—therefore cannot be a function of fit to the data. It must be a function of fit to the data plus something that biases a learner to favor one hypothesis over another, among hypotheses that fit the data equally well: this is inductive bias. Inductive bias is something that the model (or modeler) brings to the problem, not something inherent in the data.
A. Numerical data
B. Linguistic data
Data
The key is pretty. The keys are pretty.
Are the keys pretty? I like the keys to the cabinet.
The man is happy. Is the man happy?
The man who is tall is happy. I like the man who is tall.
Hypothesis 1 Hypothesis 2
{dependency}[theme=simple] {deptext} The & keys & to & the & cabinet & are …
[…]
{dependency}[theme=simple] {deptext} Is & the & man & who & is & tall & _& happy?
[…]
(a) Illustration of the role of inductive bias in generalization. A. The black points represent data, and the lines represent generalizations that could be formed on the basis of the data. Both lines fit the data exactly, but make different predictions for new datapoints. Intuitively, the blue line seems like a better hypothesis. But the choice to prefer any single generalization over any other can only be a function of inductive bias—a preference for hypotheses which are in some sense ‘simpler’—which is not a function of the data. B. Inductive bias in language. In black, possible corpus data from English. In blue, sentences which generalize from the corpus data on the basis of hierarchical structure. In red, sentences which generalize on the basis of linear word order. Intuitively, the blue sentences seem like natural generalizations and the red ones seem unnatural, even though all are equally consistent with the data given. This is claimed to be the result of innate inductive biases (Chomsky,, 1971).
Inductive bias shows up in a number of forms and under many names in the scientific literature. In linguistic theory, it appears perhaps first as the evaluation procedure in Chomsky, (1965, pp. 30–48), a function which compares two different grammars that can generate an observed set of sentences (that is, two grammars that fit some training data equally well), and ranks them in order of preference, in order to model how humans generalize beyond the data (Chomsky,, 1965, p. 45). More generally, it appears as Universal Grammar—an “innate schematism of mind that is applied to the data of experience” (Chomsky,, 1971, p. 28) to enable language learning and generalization. In this approach, (generative) grammatical theories can instantiate hypotheses about that schematism; they delimit what languages may exist and be learned from data.
This approach to linguistic theory serves the goal of explanatory adequacy: the idea is that a theory of grammar should not only capture which sentences are grammatical and ungrammatical in a particular language, but also that the theory itself should predict the generalizations that people make about language, and the languages that we actually find in the world (Chomsky,, 1965, Ch. 1). Thus if a grammatical theory seems to accommodate ‘impossible’ or unattested kinds of languages, then this would be considered evidence against that theory, because the theory is supposed to represent the “schematism of mind” or inductive bias that a learner uses to generalize from input data. This approach to explaining linguistic phenomena has been called “explanation by constrained description” (Haspelmath,, 2009, pp. 384–385).
Naturally then, inductive bias has been the major focus of study in language learning, seen as the simultaneous solution to two different problems: (1) how children can learn language from inadequate data (the classic Argument from the Poverty of the Stimulus: see Pearl,, 2022), and (2) why human language is the way it is: because if a language went against the child’s inductive biases, it would be hard or impossible for the child to learn. [^5] Given the centrality of this reasoning to much of linguistic theory, one might expect it to be backed up with strong experimental evidence that humans cannot or do not learn languages which violate the putatively universal principles of human languages. This is not the case: there is only limited and ambiguous experimental evidence for hard formal limits on human language learning. Among the most on-target existing studies is Smith et al., (1993), whose object of study was a man described as a polyglot savant living in a mental health facility. This individual and four control subjects (linguistics undergraduates) were tasked with learning artificial languages designed to be ‘impossible’ in three ways: (1) negation and tense are indicated by word order, (2) there is an agreement pattern judged to be impossible, and (3) the position of an emphatic marker is determined by a rule involving counting words. Results are not systematically reported, but seem to indicate that the polyglot was able to learn the ‘impossible’ word order and agreement rules (1) and (2), but not the rule for the emphatic marker (3). Another commonly cited experimental work on this topic is Musso et al., (2003), who expose German speakers to Italian and Japanese sentences, either following the real rules of those languages, or following modified rules deemed to be linguistically impossible, for example placing a negation marker after the third morpheme from the beginning of a sentence. The result is equally accurate learning of the ‘natural’ and ‘unnatural’ languages. fMRI on the subjects shows that the real languages elicit activity in the left inferior frontal gyrus, while the unnatural ones elicit activity elsewhere. We believe the meaning of these results is unclear. Only a small number of languages and participants (all of whom were already native speakers of largely hierarchically-structured languages) were tested, language localization in the brain is still contentious, and the patterns of brain activity for the ‘unnatural’ languages might reflect a lack of practice with such patterns, rather than their impossibility or a qualitative difference between linear and hierarchical rules. The dream has been to come up with a formalism for linguistic description which also captures human generalizations—a Universal Grammar, corresponding to a genetic endowment unique to humans and domain-specific to language, and perhaps reflecting representational structures in the brain (Chomsky,, 1988; Hauser et al.,, 2002; Berwick et al.,, 2011)—thus in one fell swoop solving (1) and (2) and also explaining why only humans have language.
Under this logic, it follows that a good theory of language learning should be restrictive: that is, there should be languages that cannot be learned under the theory, and this restriction on the hypothesis space provides explanatory adequacy. [^6] See, for example, Kodner et al.,’s (2022) criticism of Yang and Piantadosi,’s (2022) model of language learning as Bayesian program induction, a model which successfully learns grammars of various formal classes given small amounts of string input, thus addressing the Poverty of the Stimulus problem. Although the model readily meets the challenge of inducing formal structure from strings, it has been dismissed by some in the linguistics literature because the same model could also learn grammars that are unlike human language. In terms of Figure 3.2.1, the intuitively bad hypotheses would be ruled out as simply unavailable as mental representations during learning. This approach to linguistic explanation has a pleasing elegance to it: learners must be restricted to learn properly, and we see that the variation in actual languages is restricted, therefore we can kill two birds with one stone by finding the right representational restrictions on learners. [^7] For a modern and formally fleshed out take on this approach, see the Subregular Program in phonology (Heinz,, 2018) (and syntax and semantics). A group of formal linguists have argued that phonological generalizations fall into various formal classes called subregular. These formal classes have strong formal guarantees on learning from finite data when the learner is assumed to be restricted to a subregular hypothesis space.
3.2.2 The modern view on learning
The logic of inductive bias is sound. On a deep level, there really is no free lunch in learning, even deep learning (Mitchell,, 1980; Adam et al.,, 2019). However, developments in deep learning have forced massive revisions to conventional ways of thinking about inductive bias.
The conventional wisdom was that a learning model must be restricted lest it overfit (Bishop,, 2006, §1.1). That is, an unrestricted model may memorize the training data, or find some other pathological solution which lurks in the depths of its vast hypothesis space, as in the red hypotheses in Figure 3.2.1. By restricting the hypothesis space, one provides strong inductive bias, so that pathological solutions can be ruled out and good generalization can be guaranteed. The logic is the same as Universal Grammar: explaining generalization requires that we restrict the set of hypotheses (grammars) that a model (learner) may entertain.
However, the empirical finding from deep learning is that overparameterized networks, which are far more powerful than would be needed to memorize their training data, do generalize, often better than restricted ones (Belkin et al.,, 2019; Zhang et al.,, 2021). In machine learning, overfitting has typically been measured in terms of performance on a held-out dataset, where one typically sees a U-shape curve as in Figure 2A: as models become more powerful in terms of the hypotheses they can express, at first they generalize to the held-out data poorly (underfitting), then they generalize well (the sweet spot), then after some point the more powerful models learn poorer hypotheses (overfitting). And yet in the late 2010s it was discovered that if you keep adding model power, the U-shape curve starts to descend again, in a phenomenon called double descent, shown in Figure 2B, indicating that learners start to find good generalizations after being able to memorize their training data. [^8] In fact, double descent had already been discovered in the statistical physics literature (Vallet et al.,, 1989, p. 319), but its significance for machine learning was not appreciated at the time. It seems that what appeared to be an insurmountable barrier turned out to actually be a kind of phase transition in how learning works: below a certain level of model capacity, one sees the usual picture of over- and under-fitting, but above that level, one sees a new kind of generalization emerging, where more powerful learners have superior performance in learning the right generalizations.
under-fittingover-fittingsweet spotTest riskTraining riskModel capacityRiskAunder-parameterizedover-parameterizedinterpolation thresholdTest riskTraining riskModel capacityRiskB
Figure 2: Conceptual risk (probability of error) on training set and test set as a function of model capacity (the variety of hypotheses that the model can express). A. The classical picture from statistical learning theory. In the under-fitting regime, increasing model capacity reduces training and test risk, up to a point where one enters the over-fitting regime, in which increasing model capacity causes a decrease in training risk but an increase in test risk. This overfitting phenomenon is intuitively due to the model gaining the capacity to memorize the training set without forming any generalizations that would be useful on the test set. The goal of model fitting in this view is to find the sweet spot that minimizes training risk, often accomplished through methods such as deliberately reducing model capacity or early stopping. B. The modern picture arising empirically from deep learning. Classical statistical learning theory takes place in the under-parameterized regime. As model capacity increases beyond the point where the model can memorize the training data (the interpolation threshold), we enter the over-parameterized regime, where test risk decreases. Figure based on Belkin et al., (2019).
These findings came as a surprise to statistical learning theorists (Belkin et al.,, 2019; Zhang et al.,, 2021) and have triggered an ongoing field-wide effort to rethink learning theory (for example, Poggio et al., 2020a, ; Martin et al.,, 2021; Martin and Mahoney,, 2021; Kuzborskij et al.,, 2021; Henighan et al.,, 2023; Attias et al.,, 2024). We believe these results should also make us rethink language learning and the role of restrictive formalisms in linguistic explanation. It is simply not the case that proper generalization requires restricted learners, nor even that there is a correlation between restrictedness and proper generalization. All the successes of modern machine learning are in contradiction to the view that effective learners need to operate in a restricted hypothesis space.
But how can this be, if the idea of inductive bias was true? How did the conventional logic go wrong in practice? The key mistake was the conflation of model power with inductive bias, that is, the idea that inductive bias has to come from restrictions on the learner’s hypothesis space (Hubinger,, 2019). In fact, the modern empirical finding is that less restricted learners have stronger biases toward ‘simple’ hypotheses (Huh et al.,, 2024). Furthermore, the source of this simplicity bias in neural networks and related systems is not necessarily related to the hard limits of their expressivity. This is an area under active investigation—neural systems seem to be biased towards functions that are nearly linear or simple in other ways (Valle-Perez et al.,, 2018; Hahn et al., 2021b, ), likely as a result of the dynamics of gradient descent on the loss landscape induced by the model (Poggio et al., 2020b, ; Pezeshki et al.,, 2020; Merrill et al., 2021b, ; Hahn and Rofin,, 2024).
In machine learning, these discoveries have catalyzed a change in focus, away from models whose architecture and representations are tailor-fit to their domain, and toward models that learn quickly in relatively unrestricted hypothesis spaces (Sutton,, 2019). The zeitgeist is now strongly against ‘building in’ domain knowledge or targeted biases, and rather focusing on finding powerful general learners. New model architectures are now routinely evaluated based on the slope of their error rate as a function of amount of training, called scaling laws (Kaplan et al.,, 2020), rather than based on the expressivity of their architecture.
3.2.3 The upshot for linguistic theory
This revolution in learning theory means that the research program of explaining language universals through restricted learners is deprived of its apparent logical inevitability. Previously, the logic was: learners must be formally restricted, and we see that human languages do have universal patterns, so maybe these necessary restrictions create the universal patterns. Nowadays, if this approach to linguistic explanation is to be maintained, the logic must be: learners are restricted even though this is not necessary and may be harmful for learning, and these restrictions create language universals. It is still a viable hypothesis, but it no longer seems so elegant. There is no logical necessity for a restrictive formalism to explain language learning, so the approach of “explanation by constrained description” no longer seems so explanatory. Furthermore, even if evolutionarily-contingent innate inductive biases in humans really are the key to the structure of human language, these inductive biases might not be expressible in terms of a discrete formalism or a limited hypothesis space for learners.
It is still the case that, whatever language models learn about language, they do it on the basis of orders of magnitude more linguistic data than human children are exposed to (Yedetore et al.,, 2023; Warstadt et al.,, 2023). While it has been suggested that formal linguistic structure, perhaps built-in to the architecture of the model, would provide the key inductive bias to make learning from small amounts of data possible, there is no evidence that this is the only missing ingredient, or even an important one, to close the gap between models and human children. [^9] Indeed, the analysis by Mollica and Piantadosi, (2019) suggests that hierarchical syntactic structure makes up only a very small proportion of the information necessary to learn a language. It is entirely possible that the key to faster and more human-like learning is not particular built-in formal constraints, but rather more expressive architectures (such as Kolmogorov–Arnold Networks: Liu et al.,, 2024), or biases towards domain-general compositional symbolic reasoning (McCoy and Griffiths,, 2023; Yang and Piantadosi,, 2022), or simply a different training regime (Murty et al.,, 2023), or the incorporation of multimodal data which provides rich side information about the structure of the environment that is being described in language (Wang et al.,, 2023). Attempts to inject linguistic structure into neural models in various ways do succeed in promoting the kinds of structural generalizations that humans find intuitive (Dyer et al.,, 2016; Futrell et al., 2019b, ; Wilcox et al., 2019b, ; Kim et al.,, 2019), and they do succeed in making language learning from data somewhat more sample efficient (Papadimitriou and Jurafsky,, 2020, 2023), but gains in terms of absolute overall performance are small, and so such models have not caught on. Nor is it the case that formal linguists have presented an alternative model with anything like the demonstrated practical language-learning capabilities of neural models.
This is not to say that inductive biases are no longer important, nor that investigating inductive biases of humans and neural networks is not important: we will have much more to say on this in Section 4.4. Far from demoting inductive bias as a concern, language models open up the kinds of (possibly innate) inductive biases that we might look for in humans. The main upshot for linguistic theory is not to demote the importance of the learning problem, but to reshape how we think about it: away from reliance only on restrictive formalisms, and toward a broader landscape of hypotheses.
3.3 Language models and linguistic traditions
While language models may seem to involve concepts that are foreign to linguistics, they have in fact emerged in part from certain intellectual traditions of the study of language—although these traditions are not necessarily the ones that have been dominant in linguistics departments. Below we discuss the relationship between schools of linguistics and the development of language models and related language technologies.
In doing so, we focus primarily on linguistic structure, particularly related to syntax. This mirrors the historical structure of the field, and it also reflects the fact that language models make profound contributions to inquiry in these domains. But we note that linguistics also encompasses a rich tradition of the study of sociocultural aspects of language (Bucholtz and Hall,, 2005; Meyerhoff,, 2006), and that this work is increasingly important in AI. Insofar as LMs are also now (sometimes unwelcome) interlocutors in our linguistic communities due to the proliferation of widely used chatbots and the influx of automatically generated text into daily life, linguists are well-positioned to study the social and cultural role of LMs in linguistic communities.
3.3.1 The generative tradition of linguistics
It is rare in the history of science for a scientific theory to turn out as disconnected from a corresponding engineering application as formal generative linguistics has turned out to be for language models. We believe this has happened primarily because of a difference in goals between generative linguistics and language modeling, with the generative tradition taking a narrow view of its goals.
A historical parallel is informative. In the early 19th century, while Newtonian mechanics provided a strong basis for certain parts of mechanical engineering, it did not on its own seem to provide a theory that answered pressing questions about the increasingly complex machines emerging from the Industrial Revolution, in particular (1) the principles underlying steam engines, and (2) the fact that perpetual motion machines appeared to be impossible. Carnot, (1803) and then Carnot, (1824) developed a theory of these phenomena in terms of a new ad-hoc concept of ‘moment of activity’ which was dissipated inexorably by irreversible shocks to the system—irreversible at the coarse-grained level of realistic observations and average quantities, but not at the level of fine-grained mechanics which had been the focus of theoretical physics. The ‘moment of activity’ eventually developed into the idea of entropy (Clausius,, 1865) and the whole field of thermodynamics followed.
Entropy had to be discovered through engineering because the focus of purely theoretical physicists was on understanding fundamental dynamics and forces, and on developing mathematical methods of increasing elegance and sophistication. Explaining the behavior of complex machines—which behavior certainly must reduce to more basic mechanics anyway—was not on the agenda. And yet from the apparently ad-hoc project of understanding steam engines emerged a family of concepts that some theoretical physicists now view as more fundamental than even spacetime itself (Wheeler,, 1989; Cao et al.,, 2017).
Turning back to linguistics: the reason that generative linguistic theory did not turn out to be helpful along the path toward language models was due to narrowness in focus. Just as the focus of theoretical physics in the early 19th century did not lead it to consider effective theories of steam engines, the focus of generative linguistics in the 20th and early 21st centuries did not lead it to consider complex systems like language models.
Language models are, at bottom, models of the stream of language that is produced and comprehended by humans. The relevant theory is the theory of language use, production, comprehension, learning, and of human cognition more generally. Generative linguistics, on the other hand, has not focused on how to build a machine that produces and comprehends language, but rather how to build a language, conceived of as an abstract mental mapping between meaning (that is, a logical form or conceptual–intensional representation) and form (that is, a phonological form or sensorimotor representation) (Chomsky,, 2005). [^10] Early generative linguistics went further in its goals, claiming that the structures identified through generative analysis would not only provide a characterization of language, but also be a necessary and explicit part of any theory of how language could be learned and processed (Chomsky,, 1965, 1971; Bresnan,, 1982). See Stabler, (1983) for discussion of what he sees as confusion within the field as to whether or not generative linguistic theories are intended to be theories of the representations used by the brain during processing, as opposed to theories that constrain possible language. He concludes that Chomsky and others often conflated theories of grammars and theories of mental representation and processing, but that this position was not well grounded and that “it seems unlikely that linguists and psychologists really want to claim any such thing”. Furthermore, it was claimed that this kind of analysis must take center stage in the science of language, preceding any analysis of more complex systems for language processing, use, or learning (Chomsky,, 1965, Ch. 1).
Given this focus, generative linguistics has had a huge influence on the engineering field of programming language design. Indeed, the basics of all modern programming languages are based on formal linguistic theory, incorporating explicit (often context-free) grammars and parsers, as a way of linking a stream of text to a representation of a computation. Programming languages are, perhaps, exactly the ‘esoteric dialects’ spoken by ‘elaborate machines’ in the Nabokov quote that starts this section.
While this narrowness of focus was useful on its own terms, we believe it has limits. There are potentially fundamental insights to be gained now from the analysis of messy, complex, practical systems, just as happened in physics, and as is now happening in linguistics.
3.3.2 The statistical tradition of linguistics
In fact, there is a long tradition of statistical approaches to language, and these traditions were deeply involved in the early development of language models.
Perhaps the most well-known such contribution is distributional semantics: the idea that the semantics of a word is related to the distribution over contexts in which that word appears, often cited to Firth, (1957, p. 11), and developed more systematically by Harris, (1954). This idea originates out of the school of structuralist linguistics (Saussure,, 1916; Bloomfield,, 1926). Structuralist linguistics had as one theoretical aim the development of discovery procedures, which were formal procedures which could be applied to bodies of text in order to discover (and define) linguistic structures (Harris,, 1951). The most well-developed of these discovery procedures were statistical in nature. For example, Harris, (1955) developed a theory of words and morphemes based on statistical co-occurrence patterns, including a procedure for discovering morpheme boundaries by effectively calculating transitional probabilities, an idea taken up again much later in the psycholinguistics of language learning (Saffran et al.,, 1996), and closely related to tokenization methods such as byte-pair encoding (Shibata et al.,, 1999; Tanaka-Ishii,, 2021, Ch. 11). At issue was the relationship between grammatical structure and the observable statistical structure of a corpus of language—one of the core questions that linguists working on language models are interested in again today.
Harris,’s work is an intellectual precursor to modern language models. However, in practice, the statistical structuralist tradition represented by this work was largely supplanted in American linguistics departments by the generative school, which sprang from Chomsky,’s (1957) arguments that distributional statistics were irrelevant to linguistic structure and that discovery procedures were a distraction from the putatively ‘core’ questions of linguistics.
Nevertheless, the statistical analysis of language continued under the heading of usage-based approaches (Bybee and Hopper,, 2001), in which the key to language is not a set of underlying formal rules but emergent properties based on the statistics and dynamics of language use (from which rules or rule-like behavior might emerge). Traditions in typological syntax (Greenberg, 1963a, ; Dryer,, 1992), functionalist syntax (Comrie,, 1989; Keenan and Comrie,, 1977), construction grammar (Croft,, 2001; Goldberg,, 2009), probabilistic modeling (Bresnan et al.,, 2001, 2007; Christiansen and Chater,, 2016), and evolutionary linguistics (Kirby and Hurford,, 2002) have all carried the banner of a thoroughgoingly statistical approach to language.
In fact, these statistical traditions of linguistics played a catalyzing role in the birth of neural networks and LMs. While it is sometimes claimed that LMs are purely a triumph of engineering, the connectionist framework from which they emerged overlapped with the statistical tradition in linguistics, often in conflict with the anti-statistical tradition. A major locus of this early work using neural networks was about the English past tense, with Rumelhart and McClelland, (1987) developing a neural model for forming the past tense from present tense words, triggering much debate (Pinker and Prince,, 1988). Elman, 1990b, a Linguistics PhD, invented the recurrent neural network (RNN) architecture () in order to solve the problem of finding linguistic structure in time, a problem motivated by findings in linguistics and psycholinguistics (Frazier and Fodor,, 1978). This basic architecture still underlies many language models (Feng et al.,, 2024). Chris Manning, a pioneer in neural and probabilistic models, was trained as a linguist (Manning,, 1995), and argued for a statistical approach to syntax (Manning,, 2003), before going on to develop influential neural methods for performing linguistic tasks (e.g., Pennington et al.,, 2014).
The history of the field and the dominance of the generative tradition made it such that these figures are sometimes seen as ‘not linguists’. But the traditions they emerged from were fundamentally concerned with questions about human language and cognition, so we believe they do belong under the heading of linguistics. To be clear, we do not think that the success of modern language models unequivocally supports or refutes any particular intellectual tradition in linguistics. But we do think that the narrow and exclusive theoretical focus of generative linguistics, coupled with its relative dominance within American linguistics departments throughout the late 20th century, caused tragic missed connections in intellectual history.
Progress in language science will not come from hyperfocus on any one goal, nor from any one theory of what language is, nor from one framework for understanding linguistic phenomena. Rather, the science of language needs to approach as many questions as possible from as many angles as possible. In the current moment, that means leveraging the explosion of intellectual creativity springing forth around language models. Just as the concept of entropy arose from ad-hoc analysis of complex machines and ended up revolutionizing fundamental physics, it is possible and even likely that ideas based on language models will revolutionize linguistics.
4 Where does that leave the science of language?
“Trying to understand perception by studying only neurons is like trying to understand bird flight by studying only feathers: It just cannot be done. In order to understand bird flight, we have to understand aerodynamics; only then do the structure of feathers and the different shapes of birds’ wings make sense.”
Marr, (1982, p. 27)
Where do we go from here? What is the status of the formal structures discovered through linguistic analysis, if they don’t have to be innately latent in the human brain? How can we explain why language is the way it is? What approach toward language is most productive in an era when we have huge neural networks that seem to be able to produce and comprehend it?
We believe that linguistic structure is as real as it ever was, and that the LM revolution presents us with exciting new ideas and methods for studying language. Taking a functional and information-theoretic approach brings these possibilities into focus. We also consider ways in which LMs might give us reason to update some of our linguistic theories to be more gradient and usage-based, by serving as a proof-of-concept system.
From an information-theoretic perspective, as we will argue in Section 4.1, linguistic structure is real because it is an abstraction that allows us to compress the myriad data of linguistic experience in a way that lets us predict, understand, and organize that experience. Even if it is not hardwired in the brain, or anywhere, linguistic structure is a necessary concept in the sense that without it we would have no realistic hope of coherently describing linguistic data in a way that humans can understand—and that goes for language models as well as human linguistic behavior. At the same time, we think that language models can help us study what kinds of theories successfully do explain and help us understand language data and linguistic behavior.
As for answering the question of why language is the way it is, as we argue in Section 4.4, the information-theoretic perspective invites us to consider language at the functional level, where we find that human language processing and language models are more similar than one might have expected. Thus, language models can inspire explanatory theories of language in the same way that convolutional networks inspired explanatory theories of vision (see Section 3.1), and indeed we see that the neural network literature spawns multiple exciting explanatory hypotheses for human language. Considering the broader function of language—allowing humans to communicate with each other subject to constraints on how comprehension and production work—brings language models to the forefront of linguistics as indispensable tools.
4.1 Linguistic structure is real
If a general-purpose sequence model does learn language, but it does not build in the hierarchical representations common to models of linguistic structure, then what is the status of those representations and models? We believe that linguistic structure is nevertheless real in the sense of Dennett, (1989). Linguistic structure provides a compressed and useful representation of language. To explain or describe linguistic phenomena without reference to linguistic structure would be hopelessly complex, just like Marr,’s example of trying to explain bird wings by studying only feathers.
Consider a phenomenon like subject–verb number agreement. Modern language models seem to be reasonably proficient at it, even when they are trained on corpora with no examples where a verb must agree with a non-local noun (Patil et al.,, 2024). As we argued earlier, we think this happens because they are meaningfully learning grammatical structure. But consider a thought experiment: imagine that language models learned to do subject–verb agreement through some complicated and esoteric statistical process, picking up on subtle cues in data that we have never identified or thought of. Even though this complex process doesn’t seem to straightforwardly map onto our linguistic account, we find that the model is behaviorally perfect at subject–verb agreement.
Hypothetically, given that a system learned this mechanism, it might be worth investigating whether people are actually implementing some version of it, treating language models as hypothesis generators (e.g., Lakretz et al.,, 2021; Portelance and Jasbi,, 2023). If it turns out that there are in fact messy and heterogenous processes that give rise to rule-like behavior even in humans, one reaction might be to conclude that linguistic theory was wrong or misguided—that it seemed like there was such a thing as “grammatical subject”, but actually the underlying reality was more complicated and now we can do away with grammatical subject and agreement rules and replace them with the messy, complicated neural system.
This hyper-reductive conclusion would not be productive. A theory of language which points to complex neural circuits for explaining the data of subject–verb agreement, without understanding something about the concept of grammatical subject, would be like an example from Dennett, (1989, Ch. 2): imagine aliens who can perfectly predict human behavior using atomic physics (that clump of human atoms will move towards that coffee shop) but without understanding human beliefs and desires (that the human wants coffee and believes that the coffee shop will provide it). Lacking concepts of humans’ attitudes, the aliens’ theory of human behavior is perhaps perfectly accurate and reductive, but comes “at enormous cost of cumbersomeness, lack of generality, and unwanted detail” (Dennett,, 1990, p. 189). Dennett, concludes that the concept of belief is a real pattern in that it provides a (potentially not 100% accurate) abstraction that supports prediction and counterfactual reasoning based on coarse-grained data. The idea is that any abstraction that enables simple prediction, compression, and causal modeling in this way is a real pattern, fully deserving to be called real even if there is a more accurate, more reductive theory that lacks the abstraction. Similarly, even if it turned out we could explain linguistic behavior without recourse to anything recognizable as linguistic structure, we believe it would come at the same cost of cumbersomeness, lack of generality, and unwanted detail.
Figure 3: Schematic for linguistic structure as a real pattern, a (potentially leaky) abstraction that supports prediction and counterfactual reasoning based on coarse-grained data. In neural networks (in silico), the complex patterns of weights and activations may be abstracted into interpretable circuits that approximately compute linguistically meaningful features like whether a word is singular and a grammatical subject. In turn, those interpretable circuits may be interpretable in terms of a larger abstraction of a coherent linguistic structure. In humans (in vivo), the complex biological and social processes associated with language learning, use, and knowledge, may be abstracted into psycholinguistic theories describing operations on mental representations. In turn, the behaviors of these mental representations may be interpretable in terms of coherent linguistic structure. Thus linguistic structure is a real pattern at the highest level of abstraction, which allows us to understand language as it is implemented by a human or by a neural network.
Thus we join Nefdt, (2023) in seeing linguistic structure and theory as giving us real patterns—a position consistent with the position laid out in Smolensky, (1988) that “Parallel connectionist memory and linguistic processes are hypothesized to give rise to processes that are describable at a higher level as sequential rule application.” And we join Dennett, in his conviction that real patterns are, well, real. That is, we are not eliminativists about grammatical structure, nor do we believe that this kind of eliminativism would be scientifically productive. Our view is summarized in Figure 3.
Versions of our view have been a common refrain from all sides. Rumelhart and McClelland, 1986a write: “we suggest that the mechanisms that process language and make judgements of grammaticality are constructed in such a way that their performance is characterizable by rules, but that the rules themselves are not written in explicit form anywhere in the mechanism”. Some in the generative tradition seem to express the same view, like Adger, (2022): “Generative grammarians think of the posits of their theories in much the same way as physicists think of quarks or particles in wave form: They are the best explanations of the phenomena, though we do not know exactly what they correspond to in the world we can currently observe.” While there are many ways for a particular theory to be implemented computationally (Stabler,, 1983), LMs are a promising avenue for figuring out what these explanations do, in fact, correspond to at a lower level— insofar as they are successful (see earlier discussion of the Contravariance Principle) and insofar as linguistic structure provides efficient, useful representations of linguistic data.
Indeed, the real patterns framework can help explain the success of language models. It is exactly because linguistic structure is a real pattern that models are able to learn linguistic structure from data. Traditional language models are trained to predict tokens in context. From an information-theoretic perspective, this means they are finding simple representations for linguistic input (Tishby et al.,, 1999; Chater and Vitányi,, 2007; Tishby and Zaslavsky,, 2015), and in particular, compressed representations of past input that let them predict future input (Shalizi and Crutchfield,, 2001; Bialek et al.,, 2006; Still,, 2014).
Along these lines, Wolfram, (2023) makes a useful connection between LMs and the reducibility of computational processes. Some computational processes are chaotic and complex: there is no good way to systematically predict the digits of pi other than actually running the computation and learning what the next digit is. But other processes that appear very complex actually have learnable patterns that can be used to give simpler, compressed representations. For instance, $\frac{3227}{555}$ is $5.8144144144\dots$ where the $144$ repeats forever. This repeating structure means that we could give a compressed representation of what the millionth digit is without having to do additional computation. Wolfram, suggests that the success of LMs (which we know to be effective compressors) is proof that human language is amenable to this kind of compressed representation. A plausible candidate that allows for this compression is linguistic structure.
4.2 What LM experiments can in principle tell us about human language
Armed with the framework of real patterns, we can think more critically about what LM experiments can tell us about linguistic theory and why we think that it is clearly possible to learn something about human language from LM experiments. Consider a thought experiment, this time using the linguistic concept of c-command (a particular relationship between elements in a linguistic tree which has been posited to be important for a variety of structure-sensitive linguistic operations). In this thought experiment, we get embedded representations for each word and do some fancy matrix multiplication between every pair of representations for each word in each sentence, to get a representation of the pairwise relationship between each pair of words in each sentence. We do some analysis on these word pair representations, and we find that, in layer $n$, neuron $m$ patterns in a very interesting way. If and only if the value of that particular neuron is positive, then the first word in the pair c-commands the second word. If it’s negative, then we know that it doesn’t. You experiment some more and find this is remarkably and perfectly consistent. (N.B.: We want to be very clear that this is a thought experiment; no experiment would ever give such clean results and LMs just don’t seem to represent anything in such linear and human-interpretable ways.)
If we did discover a result like this, we would be justified in drawing several conclusions. First, the ease with which we extracted the c-command relationship would be incredibly compelling evidence that, for performing linguistic tasks in English, understanding which words c-command which other words is very useful exactly because, even though the model didn’t have anything about c-command built into it, it learned it emergently. Second, we would conclude that learning the c-command relationship is possible with relatively little built-in language-specific bias since a Transformer model learned it simply by optimizing for a word prediction objective. We would thus update our credence that the learning of complex syntactic relationships like c-command is possible without a built-in Universal Grammar (or that somehow the Transformer architecture is itself implementing a form of UG). Third, we could use this method to test various competing theories about which kinds of structural relationships should be more or less prominently represented in models. Perhaps most importantly, we could use these results to try to work out how an abstract structural relationship like c-command is represented in the weights of a system which, like the brain, is composed of cascades of activating neurons. Given our relatively impoverished understanding of how human neurons in the brain represent complex linguistic structures, such explorations could be informative as in principle in silico models of how the brain might process language.
It would be great if things were so easy! But they are not: the problem is that, like all complex systems, language models don’t learn and represent information in neat and interpretable ways where we can just read off what they are doing. (There is no reason for them to be interpretable; they are not optimized for it.) Rather, it takes a lot of work to ask questions like: are LMs representing part of speech (Tenney et al.,, 2019)? are they representing grammatical dependencies (Hewitt and Manning,, 2019; Manning et al.,, 2020; Chi et al.,, 2020)? how do they represent grammatical roles like subject? And the answers that come out are messy. A lot of this work has been limited by the best available techniques in interpretability. And, even if we did have perfect methods, the representations themselves would be messy and gradient, not neatly separable.
What if we looked and looked and didn’t find the c-command relationship in our model representations? That’s a trickier scenario, since we might simply not have used the right methods to find it, and it’s still robustly represented. Or it might turn out that, while c-command is critical to human processing, LMs process language in some other way that has no recourse to c-command. Or it might turn out that neither humans nor models need to represent something like c-command to demonstrate proficient linguistic behavior. Even in this negative case, though, we think the results could be fruitful for generating hypotheses and stimulating further inquiry, grounded in data and requiring precise formulations of what exactly is entailed by testing for c-command.
Thus, while the linguistics conclusions we can draw from LM intepretabilty experiments require care, we think it should be relatively uncontroversial that the idealized c-command experiment above would be linguistically interesting—even while recognizing a wide variety of opinions as to whether the current slate of results from LM experiments are linguistically interesting (Wilcox et al., 2023a, ; Mahowald et al.,, 2024; Lan et al.,, 2024).
Our own stance is that, with improvements in interpretability techniques and particularly techniques focused on generating linguistic insights into the mechanistic structure of important phenomena l, these kinds of experiments are no longer mere pipe dreams. As Chris Potts said in a talk focused on LM processing of long-distance structures (drawing on Arora et al.,, 2024): “it seems to me that I can talk about the feature in either way, symbolic or neural, without any loss of accuracy” (Potts,, 2025). This broad line of work, exploring symbolic representations in systems that solve genuinely interesting linguistic tasks, has started to make good on the promise of Smolensky, (1988)’s prescient vision “in which traditional and connectionist theoretical constructs collaborate intimately to provide an understanding of cognition.”
4.3 LMs are a proof of concept for gradient, usage-based theories of language
Under the real patterns account, we are open to the possibility that a wide variety of linguistic theories can be fruitfully combined with LMs. But the real patterns story is agnostic as to what theoretical framework will in fact be most explanatory and predictive. There is a separate question raised here, which is whether the LM era should in fact push us towards any particular theory for explaining language. We think that the success of LMs strengthens the case for existing usage-based language theories (closely related to the statistical tradition in 3.3.2) by serving as a model system for how these positions could work, and as proof of concept that they can work. By their very nature, graded usage-based theories have been difficult to formalize to the extent that one could, for example, generate or process sentences with them; language models are proofs of concept showing that the fuzzy, graded representations used in these theories can deliver a model of language.
Word Meanings
Traditional views in semantics often posit that words have discrete meanings and can be combined using systematic rules (Heim and Kratzer,, 1998). In contrast, a distributional view of meaning assumes meanings are best captured by usage patterns (Potts,, 2019; Goldberg,, 2019; Erk,, 2012; Baroni and Lenci,, 2010; Baroni et al.,, 2014; Elman, 1990a, ), an intuition now operationalized inside language models as continuous high-dimensional vectors that encode the contexts in which a word appears, resulting in a rich web of lexical relationships represented within the geometry of the vector space (Mikolov et al.,, 2013). If we think of the meaning of the word fire as being represented by its embedding in context, we find that the meaning of the word is slightly different in every context in which it occurs, but clustering together in space (campfires in a distinct cluster from a building fire even though these meanings are in some theories often lumped in as the same “meaning”; Chronis and Erk,, 2020). As such, LMs show that we might not need a strict separation between core meaning and a richly context-dependent and pragmatics-laden meaning in use. LMs give us a way of formalizing the intuition that meaning might be is gradient, fuzzy, and usage-based in a way that works in practice, not just in theory.
Syntactic Categories and Grammatical Roles
LMs seem to be able to perform tasks that require knowledge of syntactic category and grammatical role, such as generalizing part of speech when encountering a novel token (Kim and Smolensky,, 2021). One might be then tempted to ask: what are the syntactic categories an LM uses? And we can, for instance, probe for grammatical category in models and see whether the categories that emerge correspond to ones posited by linguists (Chi et al.,, 2020; Papadimitriou et al.,, 2021)?
A key finding of this work is that, while there are natural clusters in part of speech and grammatical role, such clusters are necessarily fuzzy and gradient. For instance, Papadimitriou et al., (2021) found robust representations of subjecthood but they were graded with intransitive subjects and passive subjects less “subject-y” than the subjects of transitive active verbs. These representations could be discretized to make judgments like “BERT represents this word as a subject but not this one”, but this discretization would be lossy. This graded notion of syntactic category and grammatical role has a long history in functional linguistics (Ross,, 1972) and, as such, we think can be useful in giving us a new way to think about claims that word classes are not stable across languages (Haspelmath,, 2012) and that grammatical categories exist in graded hierarchies (Comrie,, 1989; Keenan and Comrie,, 1977).
Compositionality
While traditional theories have emphasized the compositional nature of language, the construction-based approach has emphasized that not all (or even most) language can be explained strictly compositionally (Goldberg,, 2009). This position argues for a graded notion of compositionality: knowing what “green tea” is requires more than just knowing the meaning of “green” and the meaning of “tea”—even though this isn’t strictly an idiom in that the meanings of “green” and “tea” are still relevant. There has been a lot of work on whether LMs can handle strict compositionality (Lake and Baroni,, 2018; Kim and Linzen,, 2020; Russin et al.,, 2024), as well as work on how LMs handle more idiosyncratic constructions (Mahowald,, 2023; Misra and Mahowald,, 2024; Weissweiler et al., 2023a, ; Tseng et al.,, 2022; Devlin et al.,, 2019). Neural models can combine the best of both worlds in that they can maintain multiple levels of representation at once (Baroni et al.,, 2014). They don’t have to decide whether “green tea” is fully composed or fully stored: it can be a little of both. They can even be used to measure the degree of compositionality quantitatively (Rathi et al.,, 2021; Socolof et al.,, 2022).
Separation of Linguistic Layers
Traditional linguistic theory has often posited a cognitive separation in processing between linguistic layers (phonology, syntax, semantics, pragmatics). But evidence from psycholinguistics (e.g., Tanenhaus et al.,, 1995; Shain et al.,, 2024; Fedorenko et al.,, 2020) suggests these layers are not processed in neat discrete stages but are richly integrated. LMs do seem to have some separation such that lower layers of the model handle morphology with higher layers handling syntax and semantics (Tenney et al.,, 2019). But the nature of their architecture is to share information in hard-to-untangle nonlinear ways. So their functional separation is not strict but fuzzy and gradient—similar to psycholinguistic evidence for humans.
4.4 What the inductive biases of LMs can tell us about language
Since neural language models do have inductive biases, perhaps one reason for their success in learning language is because those inductive biases line up with those of humans. Indeed, a number of recent studies have demonstrated an apparent alignment of LMs’ inductive biases and the structure of human language.
In one such study, Kallini et al., (2024) refute the claim that neural LMs can learn any language, including unnatural ones, equally well (Bolhuis et al.,, 2024). They compare the learning curves for the GPT-2 architecture trained on language modeling on English text against models trained on various transformations of the English text, designed to create languages that are intuitively ‘impossible’. For example, one transformation applies a deterministic shuffling function to the tokens of English text, creating extraordinarily complex but deterministic word order rules that violate all known formal characterizations of syntax, and another transformation creates a new agreement marker that must appear 4 tokens away from a verb, also an unnatural pattern. Kallini et al., (2024) find that the model learns from real English text consistently faster than the baselines.
These results and others show that Transformers and related models do have inductive biases that align with human language in important ways. However, LM inductive biases don’t arise from restrictedness to a particular formal language class, as might be expected from the generative paradigm in linguistics. In fact, in terms of formal expressivity, it seems that Transformers are mismatched with the usual formal language classes used to characterize language. Whereas human language is sometimes characterized using (extensions of) the Chomsky–Schützenberger hierarchy (Chomsky and Schützenberger,, 1963; Weir,, 1988), which encompasses well-known classes such as regular and context-free languages, modern Transformers as they are currently applied (and other recently successful language model architectures such as State Space Models: Gu and Dao,, 2023) seem to inhabit formal language classes defined by circuit complexity (Merrill et al., 2022a, ; Strobl et al.,, 2024; Merrill et al., 2024a, ), a formal language hierarchy which is orthogonal to the Chomsky–Schützenberger hierarchy. [^11] Interestingly, circuit complexity has also been used to characterize the computational capacity of biologically realistic populations of neurons (Maass,, 1997; Maass and Markram,, 2004). So human performance may also be ultimately limited in this way. To the extent that Transformers have inductive biases that help (or hinder) their language learning, they likely arise from something other than the expressive limits of their architecture.
Below we consider some apparent learning biases of modern LMs which may be aligned with the structure of human language.
Information locality
One apparent inductive bias of autoregressive language models that seems to support language learning is information locality: a bias towards sequences where elements that statistically predict each other are close to each other. For example, in a phrase such as big brown box, the noun box and the adjective brown are highly predictive of each other—boxes, especially cardboard ones, are often brown, for many reasons—and so these words are likely to be close to each other: the alternate order brown big box sounds odd or like it is conveying some other special meaning (Futrell,, 2019; Culbertson et al.,, 2020; Scontras,, 2023; Dyer et al.,, 2023). Zooming out, paragraphs of text typically contain information about a certain topic locally, so words within a paragraph are likely to be more correlated than words spanning paragraphs. Indeed, large-scale studies of text have shown that statistical correlations among words fall off with distance (Li,, 1991; Lin and Tegmark,, 2017; Tanaka-Ishii,, 2021). Zooming in, most of the phonotactic restrictions on sound co-occurrences within words are local in nature (Heinz,, 2018).
Locality is a general property of human language (Behaghel,, 1930; Givón,, 1991): words are usually contiguous units, usually modified by prefixes and suffixes (directly adjacent to them, ordered by ‘relevance’ to the root: Bybee,, 1985; Saldana et al.,, 2024), and words linked by syntactic dependencies tend to be close to each other (Gibson,, 1991, 1998; Liu,, 2008; Liu et al.,, 2017), more than would expected under random grammars within a linguistically realistic formalism (Gildea and Temperley,, 2007; Park and Levy,, 2009; Gildea and Temperley,, 2010; Futrell et al., 2015a, ; Futrell et al., 2020b, ). These locality properties may be special cases of information locality (Futrell,, 2019; Mansfield,, 2021; Hahn et al., 2021a, ; Mansfield and Kemp,, 2023). Several of the experiments of Kallini et al., (2024) (discussed above) can be interpreted as demonstrating the information locality bias in GPT-2, because they involve baseline languages which disrupt information locality.
The bias towards locality seems to come from next-token prediction task performed by LMs. Thus, it is likely that any model with a similar task and loss function would have a similar bias. We believe that the bias toward information locality is an ‘ember of autoregression’ in the terminology of McCoy et al., (2023), one which may be shared with humans.
Relative insensitivity to word order
Transformers have a bias toward learning permutation-symmetric functions—that is, they struggle with learning functions that depend closely on the exact order of tokens within a sequence (Lavie et al.,, 2024). The effect of this bias has been observed empirically in Transformer-based language models, which early work found to be in many cases insensitive to word order (Sinha et al.,, 2021; Hessel and Schofield,, 2021; Pham et al.,, 2021; Gupta et al.,, 2021; O’Connor and Andreas,, 2021). Interestingly, humans also have difficulty remembering the exact ordering of long sequences of symbols, especially in the middle of a sequence (Henson,, 1998).
[…]
Learnability beyond inductive bias
The success of large language models invites us to go beyond thinking about learnability purely in terms of inductive bias, and to think about the limits of what any model or organism could learn, and what properties language should have in order to facilitate that learning. Addressing these issues has spurred serious new inquiry into fundamental questions about the relationship between statistical and linguistic structure, and what and how much it is possible to learn from a (large) corpus of language.
For example: Merrill et al., 2021a , drawing on Harris, (1954), state formal conditions under which (an emulation of) semantics may be acquired on the basis of distributional structure by any model, the key factor being to what extent semantics are expressed in ways that are not context-dependent—a factor deeply related to the systematic, compositional structure of human language. Merrill et al., 2022b furthermore show that entailment relations—logical relationships between propositions—can be extracted in principle from distributional statistics of text under certain assumptions about humans’ pragmatic behavior during language production, to wit that humans tend not to state the same fact more than once. The argument is formalized using the Rational Speech Acts model of pragmatics, a probabilistic game-theoretic model of language production and comprehension (Goodman and Frank,, 2016; Scontras et al.,, 2017; Degen,, 2023). Subsequently, in empirical work inspired by their theory, Merrill et al., 2024b find that entailment relations can be extracted from distributional statistics, but only under an inversion of their earlier theory—it was in fact necessary to assume a high level of redundancy in text, revising the earlier assumption.
This is just one example of a sequence of scientific models and studies, inspired by and focusing on language models, which have brought real new insight about the relationship between the structure of language (in this case, semantics), the logic of language use in context (pragmatics), and the information-theoretic structure of large bodies of language as produced by humans (distributional structure). The back-and-forth of theory development, testing against data, and theory revision in response to the data is the hallmark of healthy science.
4.5 Functional explanations for human language
Another route to explanatory adequacy avoids appeals to innateness and language-specific learning biases. The reason that ‘unnatural’ languages don’t occur may be something other than their putative innate non-learnability. Unnatural languages may not exist because they would be a poor fit to the function of language: communication and social coordination under general cognitive constraints on how language may be produced and comprehended (Chomsky,, 2005; Gibson et al.,, 2019; Levshina,, 2022; Bickel et al.,, 2024). Unsurprisingly, this function-based approach to explaining the structure of human language is associated with the functionalist school of linguistics, which is often contrasted to the generative or formalist approach (Newmeyer,, 1998), and which often overlaps with the statistical tradition discussed in Section 3.3.2. In terms of the competence–performance distinction, functionalist approaches hold that the structure of language ultimately arises from constraints and pressures on performance, which over time shape the grammars of languages that we see (Hawkins,, 1994, 2004, 2014; Haspelmath,, 2008; Comrie,, 1989).
On some level, functional explanations of human language are more basic than explanations based on learning biases. Even if we have an innate inductive bias toward human languages, this does not answer the ultimate question of why language is the way it is, it just moves the question back to why humans have one inductive bias rather than another. [^12] Indeed, a number of human learning biases tested in the artificial language learning experiments are those that also increase the communicative efficiency of language (for example, Fedzechkina et al.,, 2012; Culbertson and Adger,, 2014; Fedzechkina et al.,, 2017). From the functionalist perspective, it is no surprise that the inductive biases observed in Transformers above seem to be aligned with properties of language in key ways. They both may have their basis in deeper information-processing constraints.
One of the major findings in language comprehension has been that, like much of human cognition (Bar,, 2009; Friston and Kiebel,, 2009), it is largely predictive in nature (Ryskin and Nieuwland,, 2023): the way that processing works is influenced by the predictability of each word (or sub-word unit) conditional on previous words (Hale,, 2001; Levy,, 2008; Wilcox et al.,, 2020; Wilcox et al., 2023b, ; Xu et al.,, 2023). Furthermore, language comprehension is highly incremental (Tanenhaus et al.,, 1995; Smith and Levy,, 2013), meaning that information is processed and integrated as quickly as possible as it comes in, largely without buffering or apparent stages. There is also a high degree of incrementality in language production (Ferreira and Swets,, 2002), and incremental predictability plays a key role there too: for example, disfluencies and pauses in speech happen before unpredictable words (Goldman-Eisler,, 1957; Beattie and Butterworth,, 1979; Dammalapati et al.,, 2019), and unpredictable words are enunciated more slowly (Bell et al.,, 2009). Experimental psycholinguistics has delivered a picture of human language processing that is deeply entwined with the task of autoregressive prediction, the same task that machine learning researchers have converged on for LM training.
In general, we believe that the functional level of explanation is where linguistics and language modeling make the most fruitful contact. While bird wings and airplane wings are very different things, they share the function of flying in the Earth’s atmosphere, and so the engineering principles of aerodynamics are relevant to both airplane engineers and ornithologists (Marr,, 1982). Similarly, neural language models and human language are different things, but they share relevant principles at the level of function: they both encode information and decode information incrementally and predictively. To the extent that human language is structured to support this kind of incremental prediction (for example, through information locality), it is also aligned with the strengths and inductive biases of language models.
In fact, the explosive growth of language models and neural architectures for language provides a unique and exciting opportunity to develop the functionalist approach to language, in which models of language processing and models of language structure are like two sides of a coin (Jaeger and Tily,, 2011; Gibson et al.,, 2019; Futrell and Hahn,, 2022).
On the side of developing models of human language processing, neural language models have become an indispensable tool. Even after converging on the idea that predictive models were important for language processing, psycholinguistics as a field was hampered by the fact that strong probabilistic models for language were not available, resulting in a number of conclusions drawn on the basis of what now seem to be fairly weak models (Frank and Bod,, 2011; Fossum and Levy,, 2012; Smith and Levy,, 2013). Now, neural models make it possible to probe the limits of predictive models of language processing (Wilcox et al.,, 2020; Wilcox et al., 2023c, ; Xu et al.,, 2023), with the discovery that certain phenomena seem to go beyond the simplest predictive models (van Schijndel and Linzen,, 2018, 2021; Huang et al.,, 2024; Staub,, 2024), and that human processing is better characterized by prediction under resource constraints similar to but more severe than those present in neural models (Futrell et al., 2020a, ; Hahn et al.,, 2022; Kuribayashi et al.,, 2022; Timkey and Linzen,, 2023; Oh and Schuler,, 2023; De Varda and Marelli,, 2024).
On the side of explaining language structure, language models have again proved a key tool and source of inspiration. A sizable literature exists about what kinds of languages emerge in simulated populations of agents who communicate by encoding and decoding meanings into strings using various neural architectures, and what constraints (on the environment or the agents) are necessary for the emergent languages to resemble natural language (Lazaridou et al.,, 2017; Mordatch and Abbeel,, 2018; Steinert-Threlkeld,, 2020; Kuciński et al.,, 2021). Hahn et al., (2020) show that certain universal properties of word order, identified by Greenberg, 1963b (see also Dryer,, 1992), can be derived by finding grammars that optimize an objective function defined by (1) how easily a parse tree can be recovered from a string, and (2) how much surprisal is incurred by each word in context, with both terms operationalized using neural-network models. Because human language processing appears to be probabilistic and predictive, any theory of linguistic structure based on processing efficiency must be evaluated using a strong probabilistic predictive model, and language models provide exactly this tool. Overall, we think it is not a coincidence that the function of language models coincides so closely with the proposed function of natural language. Much of the structure of language seems to support a kind of local incremental predictability that dovetails with the function of language models (Futrell and Hahn,, 2022, 2024), and we believe that this property of language drives much of the success of autoregressive LMs.
4.6 Upshots for linguistics beyond language structure
In much of this piece, we have focused on the upshot of LMs for linguistic structure, particularly syntax. We did so for two main and related reasons: first, because of the centrality of structure for Big Questions about language over the last 60+ years; and, second, because mastery of linguistic form emerged in models earlier than other high-level abilities (Mahowald et al.,, 2024), seemingly largely from the pretraining objective. Taken together, these points have led to important insights about human language learning and processing.
But LMs have made striking gains in domains like reasoning, logic, and long-form dialog, even since the introduction of ChatGPT in 2022. These abilities seem to emerge not just from pretraining, but from techniques like supervised fine-tuning, instruction-tuning, and RLHF whereby models are given specific feedback so as to make them more aligned with human behavior on specific tasks (Bai et al.,, 2022; Achiam et al.,, 2023). The importance of inference-time compute—whereby models learn in-context and perform additional computation during the process of generating text—has also become apparent for tasks like reasoning (OpenAI et al.,, 2024). It is becoming increasingly difficult to write off these successes as mere parroting.
As such, just as it is fruitful to explore what inductive biases and representational mechanisms underlie the emergence of syntax in LMs, it is increasingly fruitful to study what inductive biases and structures are necessary and/or sufficient for other abilities in LMs like reasoning, theory of mind, planning, and other aspects of higher-level cognition. Many of our same arguments hold in these domains as well: the upshot of the modern view of learning, the contravariance principle, the role of the statistical tradition, the value of real pattern accounts.
Moreover, as LMs continue to develop, they increasingly part of our speech communities. Millions of humans regularly interact with chatbots, in some cases carrying on long extended dialogs with chatbots playing various roles. And, increasingly, text read by humans is written or co-written by AI. How will the presence of AI change language? How should we think about dialogs between human and non-human interlocutors? What social consequences will emerge from this new paradigm? Linguistics as a field, particularly sociocultural linguistics and related disciplines, is well-positioned to be at the forefront of answering these questions by drawing on an already large existing body of work and theory.
5 Conclusion
As Norm and Claudette would agree, this is a remarkable and pivotal moment in linguistics. Even ten years ago, it was not obvious that Claudette’s dream of models that produce fluent and coherent text would be possible in her lifetime—or ever. It was also not obvious that, if it were to happen, it would happen using statistical systems trained largely on next-word prediction, using massive amounts of data. These systems don’t use Norm’s hand-crafted rules. They don’t rely on insights from generative linguistic theory, although they do draw on decades of linguistic work in distributional semantics and statistical language learning.
Rather than resisting this development, we contend it offers a golden opportunity for linguistics by enabling new kinds of research and opening up vistas of new hypotheses, hypotheses that we think both Norm and Claudette will be excited about. Some of these questions will be new, like what kinds of artificial neural architectures best capture language, what kinds of biases they hold, and what the sources of those biases are. But we also have highlighted ways in which these methods can make progress on some of the oldest and most venerable questions in linguistics: what must be true of the input data for certain structures to be learned in principle, what are the constraints on what languages are possible, how does the form of language relate to its function.
Our response to all this is not to throw out linguistic structure and theory. In response to Piantadosi,’s claim that language models refute Chomsky, Kodner et al., (2023) wrote a piece skeptical of the role of LMs in linguistic theory called “Why linguistics will thrive in the 21st century”. We would go a step further and argue that, with the 21st century a quarter over, linguistics is thriving. But we think it’s doing so not in spite of language models, but in part because of them and the opportunities they open up for new research directions and the light they shed on old questions.
Therefore, we reject the dichotomy in the discourse around Piantadosi,’s (2023) paper, which often seems to pit language models against linguistics. Linguistic structure is real and important even if language models learn those structures emergently in a complex statistical way. Conversely, we think language models can point the way to new ways of thinking about the learning, processing, and fundamental nature of language.
But we do think there are some major revisions to some earlier accepted dogma that are warranted, including:
  • Much of the field of linguistics has assumed that the form of language must be explained in terms of a discrete grammar formalism that constrains the forms of possible languages, and that this grammar formalism must represent innate human constraints on language. The success of machine learning models should orient us away from this paradigm, because it shows that this approach is not necessary to characterize learning, and it opens up a new universe of statistical, quantitative, and functional theories to constrain the forms of possible languages and explain why only humans have it, while at the same time providing the tools to test those theories.
  • Relatedly, the idea that the formal structure of linguistic competence should be the main focus of theoretical linguistics, which precedes other forms of study (Chomsky,, 1965), is no longer tenable—language models reveal and allow us to characterize new dimensions of language such as how corpora of text express world knowledge, how the structure of usage supports learning or doesn’t, and how fundamental information-processing constraints shape the way that language is represented in brains and machines.
  • The success of machine learning models shows us how language can be represented in ways that are graded, probabilistic, and fuzzy, which should move us away from an insistence on discrete categorical frameworks. They also reveal potential soft constraints on the structure of language, as reviewed in Section 4.4.
  • The ontological basis for linguistic structure need not lie in an innate genetic endowment: linguistic structure is a real pattern, just as real and worthy of study whether it lies innate in the human genome or whether it is learned entirely through inductive statistical learning with domain-general biases. Linguistic phenomena do not become less interesting when they are learnable in this way, they become more interesting.
More broadly, language models show that linguistics needs the rest of science. We will not answer the most fundamental questions about language by focusing only on language, using only methods that are tailor-made for linguistic analysis. Just as neural networks revealed principles that are now fundamental in our understanding of vision, they are now in a position to revolutionize our understanding of language. Just as physics was revolutionized by thermodynamic ideas which arose from the ad-hoc analysis of complex machines, linguistics stands to benefit from new ways of thinking about computation arising from the analysis of neural networks. There is now massive intellectual activity spurred by language models in cognitive science, philosophy, physics, statistical learning theory, and information theory, and the science of language can draw ideas, methods, and inspiration from all of this by maintaining a spirit of deep, curious, open-minded engagement and integration. Human language is in many ways unique as a natural phenomenon, and this means that the science of language should integrate with other fields of science more, in order to find deeper unities and to delineate and explain how and why language is special.
In particular, we are optimistic about revising what Stabler, (1983) said in doubting that “any close connections between linguistic theory and biology will be forthcoming”. He continues: “it is a crucial advantage of the computational approach that it has a functionalist vocabulary that does not require a type reduction to physicalistic concepts […] since this seems to be required in linguistics and in psychological accounts of language processing and visual perception”. We agree that there is a crucial advantage to the functionalist vocabulary used in linguistic theory, but by analogy to recent work in vision, we are optimistic about the role that artificial neural models will play in bridging the gap between theory and biology.
Of course, a number of outstanding issues remain: there is only so much we can learn from blackbox commercial models like ChatGPT that are trained on more data than any human could ever have access to. But we can train our own smaller models on carefully controlled data and then probe them using a wide variety of techniques. While it’s true many aspects of language models were developed for engineering ends, that doesn’t mean they can’t be repurposed for scientific research. Language models are overwhelmingly models of English and other major world languages, but that too is changing as models are trained on more and more languages. These are important issues, not insurmountable roadblocks.
Baroni, (2022) lamented that, as of a 2021 exploration of citation records, Linzen et al.,’s seminal work on subject-verb agreeement was largely uncited within the linguistics community. But, since then, within the linguistics community, there are increasingly researchers using language models to ask questions within linguistics which are informed by and which can inform linguistic theory.
Similarly, while there is a persistent stereotype that the NLP community cares mostly about engineering, we have found there to be significant interest in fundamentally linguistic questions. Of the 7 papers that won Best Paper Awards at ACL 2024, 1 was a direct response to a claim by Chomsky about language learning (Kallini et al.,, 2024), 1 was about inductive biases in models and of relevance to questions about constraints on language (Hahn and Rofin,, 2024), 1 was a theoretically motivated paper studying satisfiability in natural language and of relevance to old questions about the complexity of natural language (Madusanka et al.,, 2024), 1 was a new method for measuring memorization in models and is of relevance for studying trade-offs in memorization vs. generalization in natural language (Lesci et al.,, 2024), 2 were about reconstructing or recovering ancient languages (Lu et al.,, 2024; Guan et al.,, 2024), and 1 presented an open-access mulitilingual model for broadening coverage of under-resourced languages (Üstün et al.,, 2024). Every one of these papers has relevance to scientific questions about human language, and in most cases one or more of the authors works in a linguistics department and/or has a degree in linguistics.
Thus, we don’t see ourselves as calling for radical change, but rather as acknowledging the interdisciplinary work that is already thriving in the 21st century. Linguistically informed computational work is increasingly taking place within linguistics departments, where computational researchers are working alongside syntacticians, semanticists, phonologists, language documentation experts, sociocultural linguistics, and experts in a wide variety of languages and language families. We think this is a very good development—one that both Norm and Claudette should embrace.
6 Acknowledgments
We thank Chris Potts for pointing out that he independently used a similar title in a talk called “Linguists for Deep Learning; or: How I Learned to Stop Worrying and Love Neural Networks” and encouraging us to use our title anyway. We thank Harvey Lederman for comments and helpful discussions about real patterns. We thank David Beaver, Ted Gibson, Greg Hickok, Laura Kalin, Connor Mayer, Kanishka Misra, Lisa Pearl, Greg Scontras, Steve Wechsler, and Xin Xie for helpful comments.
7 Financial Support / Funding statement
KM was supported by NSF CRII grant 2104995.
8 Competing Interest statement
The authors declare no competing interests.